OpenStudy (anonymous):

In ∆ABC, if sin A = 4/5 and tan A = 4/3 , then what is cos A? I don't understand this at all, please help.

OpenStudy (nikato):

do you have a diagram?

OpenStudy (mertsj):



OpenStudy (mertsj):

That is the triangle in which sin A= 4/5 and tan A = 4/3 What is the definition of cosine?

OpenStudy (anonymous):

All I know is that cosine is it is Adjacent / Hypotenuse.

OpenStudy (mertsj):

And, in the given triangle, what is the side adjacent to angle A?

OpenStudy (anonymous):

Angle C?

OpenStudy (mertsj):

So, your problem may be that you don't know the difference between sides and angles?

OpenStudy (anonymous):

Yeah, I got confused with the definition. Thank you for your help.

OpenStudy (mertsj):